find equation of a line. this was on my test, please solve so I can study them

A.) m=5(5,0)
B.) m=-2 (-6,2)

Answer by checkley77(12844)

A) Y=mX+b
0=5*5+b
b=-25
Y=5X-25 ANSWER FOR THIS LINE.
B) 2=-2*-6+b
2=12+b
b=2-12
b=-10
Y=-2X-10 ANSWER FOR THIS LINE.